OpenStudy (anonymous):

Let f(x) = 16x^1/3 and g(x) = 2x^2/5. Find the following: a. f(x) x g(x) b. f(x) / g(x) c. What are the domains of f x g and f / g? Thank you.

OpenStudy (solomonzelman):

a. \(\normalsize\color{royalblue}{ \rm f(x) \cdot g(x) }\) b. \(\normalsize\color{royalblue}{ \rm f(x)~/~g(x) }\) c. Domain of part a and of part b. this?

OpenStudy (anonymous):

\[f \left( x \right)=16x ^{1/3} ; g(x) = 2x ^{2/5}\] this is what f(x) and g(x) looks like.

OpenStudy (solomonzelman):

you can use ~ for a space, and if you put ~~ more space and on. (btw)

OpenStudy (solomonzelman):

and can make the size larger by putting large in front of the text

OpenStudy (anonymous):

thanks & yes those are the questions.

OpenStudy (solomonzelman):

As far as your question goes: ..... what have you done/attempted so far?

OpenStudy (solomonzelman):

do you know how to find part a? (multiply when both of the functions equal, together)

OpenStudy (solomonzelman):

\(\large\color{royalblue}{ \rm f(x) \cdot g(x)~~~~~\Rightarrow ~~~~~(16x^{{\Large 1/3}})\cdot (2x^{{\Large 2/5}})}\)

OpenStudy (solomonzelman):

can you simplify this?

OpenStudy (anonymous):

no

OpenStudy (solomonzelman):

I am having a glitch right now... I will be back as I fix it

OpenStudy (solomonzelman):

Still can't see my latex, but if you can see it, it shouldn't be a problem

OpenStudy (solomonzelman):

\(\large\color{royalblue}{ \rm 16 \cdot x^{{\Large 1/3}} \cdot 2 \cdot x^{{\Large 2/5}} }\)

OpenStudy (solomonzelman):

there, my latex is back for me

OpenStudy (solomonzelman):

\(\large\color{slate}{ \large\color{royalblue}{ \rm 16 \cdot x^{{\Large 1/3}} \cdot 2 \cdot x^{{\Large 2/5}} } }\) \(\large\color{slate}{ \large\color{royalblue}{ \rm 16 \cdot 2 \cdot x^{{\Large 1/3}} \cdot x^{{\Large 2/5}} } }\) \(\large\color{slate}{ \large\color{royalblue}{ \rm 16 \cdot 2 \cdot x^{{\Large ~1/3~+~2/5}} } }\)

OpenStudy (solomonzelman):

rewrote it for you...

OpenStudy (anonymous):

do I have to make the denominator of the fractions the same?

OpenStudy (solomonzelman):

yes.

OpenStudy (solomonzelman):

so what will 1/3 + 2/5 be? (hint: common denominator is 15)

OpenStudy (anonymous):

11/15?